Hello,
Thanks in advance for helping me out with this math problem. I am glad there is someone out there who enjoys helping students on their assignments.
The problem is as followed-
Each year in an ancient land, a large river overflowed its banks, often destroying boundary markers. The royal surveyors used a rope with knots at 12 equal intervals to help reconstruct boundaries. Explain how a surveyor could use this rope to form a right angle. (Hint: Use the Pythagorean triple 3,4, and 5)
Thank you once again....
I understand you would want to know what I think of this problem so far. I think one of the lengths of one ray of the angle is 3 and the other is 4. If it were a triangle the hypotenuse would be 5. However, once I read my answer it doesn't make sense... I know 3 times 4 equals 12? I don't know... Please help. Thank you so much

Answer by mananth(16946)

Pythagoras theorem states that side1^2+side^2=hypotenuse^2
3^2+4^2=5^2
9+16=25
25=25